Sorry, you need to enable JavaScript to visit this website.

MCST Banner Image

am1weekly_2000-134_2000-140

Mission Operations Days: 2000/134 to 2000/140
May 13, 2000 00:00:00 GMT to May 19, 2000 00:00:00 GMT


Terra Spacecraft and MODIS Instrument Status:

Terra (AM-1) is in Normal Ops Mode
MODIS is in Science Mode - A Side
MODIS has no known Anomalies
ComponentState(s)Comment
BlackbodyA On (290 K); B OffNominal
Calibration ElectronicsA On; B OffNominal
Control ProcessorA On; B OffNominal
Nadir DoorUnlatched, OpenNominal
Space View DoorUnlatched, OpenNominal
Solar Diffuser DoorUnlatched, ClosedNominal
FDDI FormatterA On; B OffNominal
FDDI PortA On; B OffNominal
FIFO Memory#1 & #2 On; #3 & #4 OffNominal
Format ProcessorA On; B OffNominal
Power Supplies#1 On, #2 OffNominal
PV VIS FPAA On; B OffNominal
PV NIR FPAA On; B OffNominal
PV SMIR FPAA On; B OffNominal
PV LWIR FPAA On; B OffNominal
PC LWIR FPAA On; B OffNominal
RC Outgas HeatersAll Off (CS, IS, OS)Nominal
RC LWIR FPA HeaterOn (83 K)Nominal
RC SMIR FPA HeaterOffNominal
Scan AssemblyA On; B OffNominal
SDSMA Off; B OffNominal
SRCAA Off; B OffNominal
PS1 Survival HeaterEnabledNominal
PS2 Survival HeaterEnabledNominal
Timing GeneratorA On; B OffNominal
Flight SoftwareRev BD + 1 set of patches
  • FR Band 14H Gain Change (10 commands)
Nominal
Inhibit Ids SetNoneNominal
TMONs Enabled66, 67Nominal

This Week’s Completed MODIS Activities:

Friday, May 12, 2000

2000/133 20:30:00 – 22:30:00 ATC OA20 10W Radiometric cont constant current w/o ND filter

Saturday, May 13, 2000

2000/134 02:30:00 – 06:10:00 ATC OA21 1W Radiometric cont constant current for 2 orbits
2000/134 10:49:58 ATC Set SciAbnormal Flag
2000/134 10:50:00 – 11:22:33 ATC OA25 1W Spatial w/ sector rotation
2000/134 11:22:35 ATC Clear SciAbnormal Flag
2000/134 15:49:58 ATC Set SciAbnormal Flag
2000/134 15:50:00 – 16:22:33 ATC OA25 10W Spatial w/slit reticle and sector rotation
2000/134 16:22:35 ATC Clear SciAbnormal Flag
2000/134 17:29:21 - 17:51:17 DATA LOSS - Unrecoverable

Sunday, May 14, 2000

None

Monday, May 15, 2000

2000/136 12:00:00 ATC Blackbody to 270K

Tuesday, May 16, 2000

2000/137 06:00:00 ATC Blackbody to 280K
2000/137 10:00:00 ATC Blackbody to 285K
2000/137 12:00:00 ATC Blackbody to 290K
2000/137 13:57:35 Real-time Set Blackbody duty cycle to FULL
2000/137 14:00:00 ATC Blackbody to 295K
2000/137 16:00:00 ATC Blackbody to 300K
2000/137 18:00:00 ATC Blackbody to 315K
2000/137 20:00:00 ATC Blackbody to 270K
2000/137 20:00:00 Real-time Set Blackbody duty cycle to THIRD

Wednesday, May 17, 2000

2000/138 19:00:00 ATC Blackbody to 290K
2000/138 19:10:00 Drag make-up maneuver - successful

Thursday, May 18, 2000

None

Friday, May 19, 2000

None

This Week's Scheduled MODIS Activities Not Completed:

No SD/SDSM activities were run this week. Monday's activities were deleted from the timeline during the system upgrade. A DR was submitted for this anomaly. Wednesday's activities were not scheduled due to the drag make-up burn which occurred that day. Friday's activities were not scheduled because of planning aid issues related to the burn. The nominal SD/SDSM activity schedule should resume next week.

Upcoming MODIS Events:

Saturday, May 20, 2000

None

Sunday, May 21, 2000

None

Monday, May 22, 2000

2000/143 ?? ATC - OA15 SD/SDSM Open
2000/143 ?? ATC - OA16 SD/SDSM Screened
2000/143 20:02:00 ATC - Blackbody to 270K

Tuesday, May 23, 2000

2000/144 14:00:00 ATC Blackbody to 315K
2000/144 16:22 Real-time Set Blackbody duty cycle to 100%
2000/144 21:00:00 ATC Blackbody to 270K
2000/144 21:23 Real-time Set Blackbody duty cycle to 33%

Wednesday, May 24, 2000

2000/145 19:00:00 ATC Blackbody to 290K
2000/145 ?? ATC OA16 SD/SDSM Screened

Thursday, May 25, 2000

None

Friday, May 26, 2000

2000/147 ?? ATC OA16 SD/SDSM Screened

Saturday, May 27, 2000

None

Sunday, May 28, 2000

None

Maneuvers:

Wed, June 21, 2000 – MODIS Roll (2000/173/23:28:43 – 23:35:13), roll degree = -17.45

MODIS Anomalies:

None.

General Instrument Comments:

MODIS is in Science Mode on the A-side with the SVD and NAD open.

MODIS Telemetry Trends:

Telemetry is nominal.

Non-MODIS Significant Events:

The Ampex tape drive issue at White Sands has been resolved. Nominal data flow resumed on Day 139/00:00:00. Backlog reprocessing continues both at EDOS and the DAAC.

Limited Life Item Status:

SRCA 10W Lamp #1: 163.9 of 500 hours
SRCA 10W Lamp #2: 131.0 of 500 hours
SRCA 10W Lamp #3: 140.2 of 500 hours
SRCA 10W Lamp #4: 61.5 of 500 hours

SRCA 1W Lamp #1: 554.3 of 4000 hours
SRCA 1W Lamp #2: 276.3 of 4000 hours

Solar Diffuser Door: 1343 of 3022 Movements
Nadir Aperture Door: 532 of 1316 Movements
Space View Door: 437 of 1316 Movements